OpenStudy (anonymous):

Find a point-slope form for the line that satisfies the conditions given. When two points are given, use the first point in the point-slope form.

OpenStudy (anonymous):

let the point be(x1,y1) by problem, (y-y1)=m(x-x1) where m=slope of the line x1,y1 is a point

OpenStudy (anonymous):

I have an x intercept of 1/2 and a y intercept of -1